Home Assignment 1

In the table below is a list of accidents that occurred in different chemical plants. Do a literature search on one of these incidents and write a report the focusses on the following aspects:

1. Describe the incident.

2. Describe the specific chemical process (include a flow diagram of the process).

3. Describe how the process was supposed to behave at steady state and which variables were not at steady state as a result of the incident.

Bhopal, India  1984
Basel, Switzerland  1986
Bradford, United Kingdom  1992
Castleford, United Kingdom 1992
Feyzin, France  1966
Flixborough, United Kingdom 1974
Guadalajara, Mexico  1992
Mexico City, Mexico  1984
Seveso, Italy  1976
Texas City, USA  1947

This is a group assignment. Please submit only one copy of the report per group.